# EntroPE: Entropy-Guided Dynamic Patch Encoder for Time Series Forecasting

### v1

@@ 3.3 Entropy-Based Dynamic Patcher (EDP)
Drawing inspiration from the decoder-only transformer architecture and cross-entropy training demonstrated in Chronos, we adopt the GPT-2 paradigm (Radford et al., 2019) to train a compact transformer for next-token prediction on quantized time series tokens. Our lightweight architecture comprises 8 embedding dimensions (shared input and head), 2 transformer layers, 4 attention heads, totaling 3,648 learnable parameters. The model performs autoregressive prediction at time $\{t_{2},t_{3},\ldots,t_{L}\}$ from input sequence at time $\{t_{1},t_{2},\ldots,t_{L-1}\}$ using standard cross-entropy loss.
This streamlined design efficiently captures temporal dynamics while maintaining computational tractability for entropy estimation. Crucially, our objective is not to achieve superior forecasting performance, but rather to align the causal transformer for measuring predictive uncertainty across the discrete vocabulary. This architectural approach enables identification of natural transition points in raw time series sequences, where predictive uncertainty that quantified through entropy over the output next-token distribution reaches local maxima.
Entropy-Guided Patch Boundary Detection. Upon training convergence, we freeze model weights and compute entropy for each time position. For position $t$, we calculate the Shannon entropy of the next-token prediction distribution:
$H(x_{t})=-\sum_{v\in\mathcal{V}}p_{\theta}(x_{t+1}=v|x_{\leq t})\log p_{\theta}(x_{t+1}=v|x_{\leq t})$ (1)
Following Pagnoni et al. (2025), we implement dual-threshold boundary detection combining global and relative uncertainty measures. A patch boundary is placed at position $t$ when three conditions are met simultaneously:
$H(x_{t})$ $>\theta\quad\text{(global entropy threshold)}$ (2)
$H(x_{t})-H(x_{t-1})$ $>\gamma\quad\text{(relative entropy increase)}$ (3)
This mechanism ensures patches begin only at positions with both high absolute uncertainty and significant uncertainty increases, avoiding fragmentation from minor entropy fluctuations while capturing meaningful temporal transitions. EDP outputs patch mask $M$, which indicates the non-overlapping patch segmentation for further processing.
@@ 3.4 Adaptive Patch Encoder (APE)
The EDP produces variable-length patches where both patch count $\hat{p}$ and individual lengths vary across samples (patch mask $M$). As shown in Fig. 2(B), the APE converts these heterogeneous patches into fixed-size representations suitable for transformer processing while preserving temporal information. When handling variable patches, we ensure batch processing compatibility by padding sequences to the maximum patch count within each batch and applying attention masks to handle variable patch lengths during cross-attention operations.
Architecture Design. Our encoder employs a two-stage approach. First initial dimensionality reduction via pooling (from time-point embeddings to initial patch embeddings), followed by iterative cross-attention refinement. This design is inspired by the Perceiver architecture (Jaegle et al., 2021) and Byte Latent Transformer (Pagnoni et al., 2025), but specialized for temporal patch aggregation.
Initial Patch Embedding. We first obtain the time-point embedding $h_{0}$ for the entire sequence with length $L$ through an Embedding Layer $E\in\mathbb{R}^{V\times d_{t}}$ over the same Vocabulary and Tokenizer used in the Entropy Model. For each variable-length patch $p_{j}$, we generate initial patch representations via max pooling: $P_{0,j}=\text{MaxPool}(E_{f}(p_{j}))$, where $E_{f}$ represents the time-point embedding function. Max pooling captures salient temporal features while providing translation invariance within patches.
Cross-Attention Refinement. We refine patch representations using cross-attention, where patch embeddings serve as queries and time-point embeddings act as keys and values. For layer $n$:
$P_{n}=P_{n-1}+W_{o}\left(\text{softmax}\!\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V\right),$ (4)
with attention components:
$Q_{j}$ $=W_{q}(P_{n-1,j})\quad\text{(patch embeddings as queries)},$ (5)
$K_{i}$ $=W_{k}(h_{n-1,i}),\quad V_{i}=W_{v}(h_{n-1,i})\quad\text{(time-point embeddings as keys/values)},$ (6)
where $W_{q}$, $W_{k}$, $W_{v}$, and $W_{o}$ are learnable projection matrices corresponding to queries, keys, values, and output transformations, respectively. For $N$-layer architectures with $N>1$, time-point embeddings are iteratively refined through stacked transformer layers as $h_{n}=\text{APETransformer}(h_{n-1})$, with cross-attention mechanisms operating between consecutive layers. Here, $h_{n}$ is the time point embedding sequence at layer $n$.
To ensure that self-attention within each APETransformer operates exclusively within patch boundaries, we employ a tailored attention mask that accommodates the variable patch lengths throughout the sequence. This masking strategy prevents information leakage across patch boundaries while preserving intra-patch temporal dependencies.
The encoder outputs fixed-size patch embeddings $P^{in}=P_{N}\in\mathbb{R}^{B\cdot C\times\hat{p}\times d_{p}}$ for global transformer processing and preserves encoder hidden states $H^{enc}=h_{N}\in\mathbb{R}^{B\cdot C\times L\times d_{t}}$ for temporal information in the decoding stage.
@@ 3.5 Global Transformer & Fusion Decoder
The final components of the EntroPE process are the fixed-size patch embeddings to learn long-term dependencies and generate forecasts.
Global Transformer. The patch embeddings $P^{in}$ from the adaptive encoder are processed through a standard transformer architecture to capture long-range temporal relationships between patches. Since intra-patch dependencies are already modeled by the encoder, the global transformer focuses exclusively on inter-patch interactions. We employ bidirectional attention (non-causal) to enable comprehensive context modeling across the entire sequence: $P^{out}=\text{GlobalTransformer}(P^{in})$.
Fusion Decoder (FD). A key challenge in dynamic patching is that the number of patches $\hat{p}$ varies across samples, making direct forecasting from flattened patch representations problematic. To address this, we employ a cross-attention mechanism that fuses global patch context with fine-grained temporal information preserved from the encoder (see Fig. 2(C)). The decoder reverses the attention configuration used in the adaptive encoder. Here, encoder hidden states $H^{enc}$ serve as queries, while global patch embeddings act as keys and values:
$Q_{i}$ $=W_{q}(H^{enc}_{i})\quad\text{(time-point queries)},$ (7)
$K_{j}$ $=W_{k}(P^{out}_{j}),\quad V_{j}=W_{v}(P^{out}_{j})\quad\text{(patch keys/values)}$ (8)
@@
The cross-attention operation enriches time-point representations with global context:
@@
$H^{dec}=H^{enc}+W_{o}\left(\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V\right).$ (9)
@@
This mechanism enables knowledge transfer from high-level patch representations back to detailed time-point embeddings, ensuring preservation of both local temporal patterns and global contextual dependencies. Optionally, similar to APE, the iterative cross-attention refinement can be configured through alternating transformer layers and cross-attention layers, creating a structured information exchange between different representational levels.

### v2

@@ 3.3 Entropy Model Pre-training
Model Selection. Drawing inspiration from decoder-only transformers in time series (Das et al., 2024) and cross-entropy guided training in Chronos, we adopt the GPT-2 paradigm (Radford et al., 2019) to train a compact transformer for next-token prediction on quantized time series tokens. Our lightweight architecture comprises 8-16 embedding dimensions and 2 transformer layers, totaling $\approx 10$k learnable parameters. The model performs autoregressive prediction at timesteps $\{t_{2},t_{3},\ldots,t_{L}\}$ from input sequence $\{t_{1},t_{2},\ldots,t_{L-1}\}$.
Model Training. We train our lightweight causal transformer $f_{\theta_{\text{ent}}}:\mathcal{V}^{t}\to\Delta^{V}$ (where $\Delta^{V}$ denotes the probability simplex over vocabulary $\mathcal{V}$) to perform next-token prediction. The model is optimized via:
$\theta_{\text{ent}}^{*}=\arg\min_{\theta_{\text{ent}}}\mathbb{E}_{X\sim\mathcal{D}_{\text{pretrain}}}\left[\mathcal{L}_{\text{CE}}(X;\theta_{\text{ent}})\right]$ (2)
where $\mathcal{L}_{\text{CE}}(X;\theta_{\text{ent}})$ is the cross-entropy loss (Eq. 21).
Once converged, we freeze $\theta_{\text{ent}}^{*}$ and use it exclusively for entropy computation, not for prediction or token generation. This decoupling is critical, as the entropy model learns predictive uncertainty patterns, while downstream task models operate on continuous representations.
3.4 Entropy-Guided Dynamic Patcher (EDP)
Entropy Computation.
For a new input sequence $X\in\mathbb{R}^{C\times L}$, we process each channel independently. For channel $c$, we compute the quantized sequence $\tau^{(c)}=[q(x_{1}^{(c)}),\ldots,q(x_{L}^{(c)})]$ and use the frozen model $\theta_{\text{ent}}^{*}$ to calculate conditional entropy at each position:
$H(x_{t}^{(c)};\theta_{\text{ent}}^{*})=\\ -\sum_{v\in\mathcal{V}}p_{\theta_{\text{ent}}^{*}}(\tau_{t+1}^{(c)}=v|\tau_{\leq t}^{(c)})\log p_{\theta_{\text{ent}}^{*}}(\tau_{t+1}^{(c)}=v|\tau_{\leq t}^{(c)})$ (3)
yielding entropy sequence,
$\mathbf{H}^{(c)}=[H(x_{1}^{(c)}),\ldots,H(x_{L}^{(c)})]\in\mathbb{R}^{L}.$ (4)
As established in Appendix A.2, high conditional entropy indicates positions where the future is less predictable from the past, typically at regime shifts, trend reversals, or volatility bursts. By placing boundaries at such points (Proposition A.2), patches become internally coherent while transition uncertainty is localized at edges.
Adaptive Threshold Determination.
Rather than using fixed global thresholds requiring manual tuning, we employ sample-adaptive thresholds derived from quantile-based estimation. For each input sequence with entropies $\{H(x_{1}^{(c)}),\ldots,H(x_{L}^{(c)})\}$, we compute sample-specific thresholds. Hence, the threshold selection becomes interpretable.
The absolute entropy threshold is the $\alpha$-th quantile of the sequence’s entropy distribution:
$\theta_{\text{abs}}^{(c)}=Q_{\alpha}(\mathbf{H}^{(c)}).$ (5)
The relative threshold is the $\alpha$-th quantile of consecutive differences $\Delta\mathbf{H}^{(c)}=\bigl[H(x_{t}^{(c)})-H(x_{t-1}^{(c)})\bigr]_{t=2}^{L}$,
$\gamma_{\mathrm{rel}}^{(c)}$ $=Q_{\alpha}\!\left(\Delta\mathbf{H}^{(c)}\right),$ (6)
where $\alpha$ is the quantile parameter (e.g., $\alpha=75\%$). This formulation enables automatic adaptation to varying entropy levels across sequences, eliminating dataset-specific threshold tuning.
Boundary Detection.
A patch boundary is placed at position $t$ based on the following conditions:
$H(x_{t}^{(c)})$ $>\theta_{\text{abs}}^{(c)}\quad\text{(high absolute uncertainty)}$ (7)
$\Delta H(x_{t}^{(c)})$ $>\gamma_{\text{rel}}^{(c)}\quad\text{(significant uncertainty increase)}$ (8)
$M_{t-1}^{(c)}$ $=0\quad\text{(no consecutive boundaries)}.$ (9)
This threshold mechanism produces a binary patch boundary mask $M^{(c)}\in\{0,1\}^{L}$ where $M_{t}^{(c)}=1$ indicates a patch start. The framework supports using either threshold independently or in combination, allowing flexible adaptation to different temporal characteristics. EDP outputs mask $M$ for subsequent non-overlapping patch processing.
Patch Segmentation.
Given boundary mask $M^{(c)}$, we extract boundary positions $\mathcal{B}^{(c)}=\{t:M_{t}^{(c)}=1\}=\{b_{1}^{(c)},b_{2}^{(c)},\ldots,b_{P^{(c)}+1}^{(c)}\}$ where $b_{1}^{(c)}=1$ and $b_{P^{(c)}+1}^{(c)}=L+1$. The sequence is then partitioned into $P^{(c)}$ non-overlapping patches:
$x^{(c)}=\bigcup_{j=1}^{P^{(c)}}p_{j}^{(c)},$ (10)
where $\quad p_{j}^{(c)}=[x_{b_{j}^{(c)}},x_{b_{j}^{(c)}+1},\ldots,x_{b_{j+1}^{(c)}-1}]\in\mathbb{R}^{\ell_{j}^{(c)}}$, $\ell_{j}^{(c)}=b_{j+1}^{(c)}-b_{j}^{(c)}$ is the length of patch $j$, and $\sum_{j=1}^{P^{(c)}}\ell_{j}^{(c)}=L$. Our segmentation depends only on relative uncertainty structure, not absolute probability calibration.
@@ 3.5 Adaptive Patch Encoder (APE)
The EDP produces variable-length patches where both patch count and individual lengths vary across samples. As shown in Fig. 2(B), the APE converts these heterogeneous patches into fixed-size representations suitable for transformer processing while preserving temporal information. When handling variable patches, we ensure batch processing compatibility by padding sequences to the maximum patch count within each batch and applying attention masks to handle variable patch lengths during cross-attention operations.
Architecture Design. Our encoder employs a two-stage approach. First initial dimensionality reduction via pooling (from time-point embeddings to initial patch embeddings), followed by iterative cross-attention refinement. This design choice is inspired by the Perceiver architecture (Jaegle et al., 2021) and Byte Latent Transformer (Pagnoni et al., 2025), due to their demonstrated performance on language and image data.
Patch Encoding.
Each variable-length patch $p_{j}^{(c)}$ is encoded into a fixed-dimensional representation via the APE. First, we obtain time-point embeddings through an embedding layer $E:\mathcal{V}\to\mathbb{R}^{d_{t}}$:
$\mathbf{h}_{t}^{(c)}=E(q(x_{t}^{(c)})),\quad t=1,\ldots,L.$ (11)
For patch $p_{j}^{(c)}$ spanning positions $\mathcal{T}_{j}^{(c)}=\{b_{j}^{(c)},b_{j}^{(c)}+1,\ldots,b_{j+1}^{(c)}-1\}$, we initialize its representation via max pooling:
$\mathbf{z}_{j}^{(c),0}=\text{MaxPool}\left(\{\mathbf{h}_{t}^{(c)}\}_{t\in\mathcal{T}_{j}^{(c)}}\right)\in\mathbb{R}^{d_{p}}.$ (12)
We then refine patch embeddings through $N$ cross-attention layers where patch embeddings query only their constituent time points:
$\mathbf{z}_{j}^{(c),n}=\mathbf{z}_{j}^{(c),n-1}+W_{o}\sum_{i\in\mathcal{T}_{j}^{(c)}}\alpha_{ji}^{(n)}W_{v}(\mathbf{h}_{i}^{(c),n-1})$ (13)
where attention weights are computed only within the patch:
$\alpha_{ji}^{(n)}=\frac{\exp\left(\frac{[W_{q}(\mathbf{z}_{j}^{(c),n-1})]^{\top}W_{k}(\mathbf{h}_{i}^{(c),n-1})}{\sqrt{d_{k}}}\right)}{\sum_{i^{\prime}\in\mathcal{T}_{j}^{(c)}}\exp\left(\frac{[W_{q}(\mathbf{z}_{j}^{(c),n-1})]^{\top}W_{k}(\mathbf{h}_{i^{\prime}}^{(c),n-1})}{\sqrt{d_{k}}}\right)}$ (14)
This produces final patch embeddings $P^{(c)}=[\mathbf{z}_{1}^{(c),N},\ldots,\mathbf{z}_{P^{(c)}}^{(c),N}]\in\mathbb{R}^{P^{(c)}\times d_{p}}$ and encoder hidden states $H^{(c)}=[\mathbf{h}_{1}^{(c),N},\ldots,\mathbf{h}_{L}^{(c),N}]\in\mathbb{R}^{L\times d_{t}}$.
Batch Processing with Variable Patches.
For a batch of $B$ samples with varying patch counts $\{P^{(b,c)}\}$, let $P_{\max}=\max_{b,c}P^{(b,c)}$. We construct a padded patch embedding tensor $\tilde{P}^{in}\in\mathbb{R}^{(B\cdot C)\times P_{\max}\times d_{p}}$ with attention mask $\mathcal{M}\in\{0,1\}^{(B\cdot C)\times P_{\max}}$ where $\mathcal{M}_{bc,j}=\mathbb{1}[j\leq P^{(b,c)}]$ to mask padded positions during self-attention.
@@ 3.6 Global Transformer & Fusion Decoder
The final components of the EntroPE process are the fixed-size patch embeddings to learn long-term dependencies and generate predictions.
Global Transformer.
Patch embeddings are processed through a standard transformer to capture inter-patch dependencies:
$\tilde{P}^{(c)}=\text{Transformer}(P^{(c)},\mathcal{M})\in\mathbb{R}^{P^{(c)}\times d_{p}}$ (15)
Fusion Decoder.
@@
We fuse global patch context with fine-grained temporal information via cross-attention where time-point embeddings query patch representations:
@@
$\tilde{H}^{(c)}=H^{(c)}+\\ W_{o}\left(\text{softmax}\left(\frac{[W_{q}(H^{(c)})][W_{k}(\tilde{P}^{(c)})]^{\top}}{\sqrt{d_{k}}}\right)W_{v}(\tilde{P}^{(c)})\right)\\ \in\mathbb{R}^{L\times d_{t}}$ (16)
@@
This mechanism enables knowledge transfer from high-level patch representations back to detailed time-point embeddings, preserving both local temporal patterns and global contextual dependencies.
Task-Specific Head. The enriched representations $\tilde{H}^{(c)}$ are projected to generate future predictions. We apply flattening followed by a linear projection with Instance De-Normalization: $\text{InstanceDeNorm}\!\left(W_{\text{proj}}(\text{Flatten}(\tilde{H}^{(c)}))\right)$. Training criteria and additional details on task-specific heads are provided in Appendix B.
4 Experiments
In the section we evaluate the efficacy of EntroPE on long-term forecasting, classification and anomaly detection tasks. We show that our model can serve as a foundation model with competitive performance accross these tasks.
4.1 Long-Term Forecasting
Datasets. To examine forecasting performance we conduct comprehensive evaluations on six widely-used long-term multivariate forecasting datasets. i.e., Electricity (ECL) featuring electricity consumption data, ETT family (ETTh1, ETTh2, ETTm1, ETTm2) that encompass a range of scenarios in energy transfer technology, Weather that offers insights into various meteorological variables over time. Detailed descriptions in Appendix Tab. 6.
Baselines and Experimental Settings. We benchmark against fifteen state-of-the-art baselines. (1) Transformer-based models: TimeMosaic (Ding et al., 2025), iTransformer (Liu et al., 2024b), PatchTST (Nie et al., 2023), FEDformer (Zhou et al., 2022), Autoformer (Wu et al., 2021); (2) CNN-based model: TimesNet (Wu et al., 2023), TSLANet (Eldele et al., 2024); (3) MLP-based models: TimeMixer (Wang et al., 2024), HDMixer (Huang et al., 2024), DLinear (Zeng et al., 2023); (4) Foundation and lightweight models: Time-FFM (Liu et al., 2024a), TimeBase (Huang et al., 2025a); (5) LLM-aligned and cross-modal models: LangTime (Niu et al., 2025), CALF (Liu et al., 2025b); (6) Filter-based model: FilterTS (Wang et al., 2025). (Further details on baselines are discussed in Appendix D.1.)
Table 1: Multivariate time series forecasting results on benchmark datasets. Results are averaged across prediction horizons $T=\{96,192,336,720\}$ with fixed input length $L=96$ for all datasets. Best results are highlighted in red and second/-best in blue.
Models ETTh1 ETTh2 ETTm1 ETTm2 Weather Electricity
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Autoformer [2021] 0.496 0.487 0.450 0.459 0.588 0.517 0.327 0.371 0.338 0.382 0.227 0.338
FEDformer [2022] 0.498 0.484 0.437 0.449 0.448 0.452 0.305 0.349 0.309 0.360 0.214 0.327
DLinear [2023] 0.461 0.457 0.563 0.519 0.404 0.408 0.354 0.402 0.265 0.315 0.225 0.319
TimesNet [2023] 0.495 0.450 0.414 0.427 0.400 0.406 0.291 0.333 0.251 0.294 0.193 0.304
PatchTST [2023] 0.516 0.484 0.391 0.411 0.406 0.407 0.290 0.334 0.265 0.285 0.216 0.318
Time-FFM [2024] 0.442 0.434 0.382 0.406 0.399 0.402 0.286 0.332 0.270 0.288 0.216 0.299
HDMixer [2024] 0.448 0.437 0.384 0.407 0.396 0.402 0.286 0.331 0.253 0.285 0.205 0.295
iTransformer [2024] 0.454 0.447 0.383 0.407 0.407 0.410 0.288 0.332 0.258 0.278 0.178 0.270
TimeMixer [2024] 0.459 0.444 0.390 0.409 0.382 0.397 0.279 0.324 0.245 0.276 0.182 0.272
TSLANet [2024] 0.448 0.441 0.372 0.399 0.378 0.397 0.283 0.327 0.259 0.280 0.199 0.283
TimeBase [2025] 0.463 0.429 0.409 0.425 0.431 0.420 0.290 0.332 0.252 0.279 0.227 0.296
LangTime [2025] 0.437 0.425 0.375 0.392 0.397 0.392 0.284 0.321 0.252 0.273 0.201 0.285
FilterTS [2025] 0.440 0.432 0.375 0.399 0.386 0.397 0.279 0.323 0.253 0.280 0.184 0.275
CALF [2025] 0.441 0.435 0.372 0.395 0.396 0.391 0.280 0.321 0.250 0.274 0.177 0.266
TimeMosaic [2026] 0.427 0.442 0.369 0.399 0.388 0.392 0.278 0.319 0.253 0.277 0.188 0.281
EntroPE [2026] 0.416 0.425 0.366 0.387 0.378 0.391 0.286 0.335 0.242 0.273 0.182 0.271
Results.
Table 1 shows EntroPE achieves competitive or superior performance across all baseline categories (Transformer, CNN, MLP, foundation, LLM-aligned, and filter-based models). While retaining PatchTST-like architecture, EntroPE introduces dynamic patching and adaptive encoding, yielding  20% MSE reduction on ETTh1,  15% on Electricity, and  10% average improvement over PatchTST, alongside reduced token count through non-overlapping patches. Notably, iTransformer and CALF demonstrates remarkable performance on high-dimensional Electricity dataset (321 variables) due to its channel-wise attention design. While CALF ($\sim$18M parameters) and LangTime ($\sim$500M parameters) show competitive results, EntroPE achieves comparable accuracy with 500-1000× fewer parameters (100k-1M), demonstrating that entropy-guided segmentation provides robust improvements without requiring massive model capacity.
Compared to some noteworthy dynamic or adaptive patching approaches such as HDMixer and TimeMosaic, EntroPE consistently delivers stronger or more stable gains across datasets. Notably, TimeMosaic employs substantially larger models (3–13M parameters) and higher training durations, whereas EntroPE achieves comparable or better accuracy with smaller models. This highlights that performance gains arise from principled entropy-driven boundary placement rather than increased model capacity. (Full table: Tab. 16 & 17)
4.2 Classification
Datasets. To demonstrate the generalizability of EntroPE beyond forecasting, we evaluate on 10 UEA multivariate time series classification datasets (Bagnall et al., 2018). We replace the forecasting head with a classification head while keeping the core architecture unchanged.
Baselines and Experimental Settings.
We select eight state-of-the-art baselines, i.e., TSLANet, GPT4TS (Zhou et al., 2023), TimesNet, ROCKET (Dempster et al., 2020), TS-TCC (Eldele et al., 2021), Crossformer (Zhang and Yan, 2023) and PatchTST as they showed the best classification accuracy over other Transformer-based architectures. Last, we experiment with a simple single-layer MLP.
Table 2: Classification accuracy (%) on 10 UEA multivariate classification datasets.
Dataset EntroPE TSLANet GPT4TS TimesNet ROCKET CrossF. PatchTST MLP TS-TCC
[2026] [2024] [2023] [2023] [2020] [2023] [2023] [-] [2021]
EthanolConcentration 29.1 30.42 25.48 27.73 42.58 34.98 28.90 33.46 32.32
FaceDetection 69.1 66.77 65.58 67.47 64.70 66.17 68.96 67.42 63.05
Handwriting 57.4 57.88 34.56 26.18 48.47 26.24 26.00 22.47 47.76
Heartbeat 77.9 77.56 36.59 74.48 69.76 76.59 76.59 73.17 77.07
JapaneseVowels 99.2 99.19 98.11 97.83 95.68 98.92 98.65 97.84 97.30
PEMS-SF 88.6 83.82 87.28 88.13 75.10 82.08 88.44 82.08 86.71
SelfRegulationSCP1 92.9 91.81 91.47 77.43 84.64 92.49 89.76 88.40 91.13
SelfRegulationSCP2 63.1 61.67 51.67 52.84 54.44 53.33 54.44 51.67 53.89
SpokenArabicDigits 99.8 99.91 99.36 98.36 99.20 96.41 99.68 96.68 99.77
UWaveGestureLibrary 94.4 91.25 84.38 83.13 76.03 81.56 80.00 81.88 86.25
Average 77.13 76.03 67.45 69.36 72.90 70.88 71.14 69.51 73.53
Results.
Table 2 shows EntroPE achieves an average accuracy of 77.13%, competitive with TSLANet (76.03%), and outperforming TimesNet (69.36%), PatchTST (71.14%), and other competitive baselines. This demonstrates that entropy-guided dynamic patching effectively captures discriminative temporal patterns for classification tasks.
Anomaly Detection Evaluation.
Appendix C shows the detailed information on datasets, baselines and results of EntroPE’s Anomaly detection performance.
4.3 Ablation Study
Model component analysis. To evaluate each component’s contribution, we conduct comprehensive ablation studies across four datasets (ETTh1, ETTh2, ETTm1, Weather) using prediction horizons of 336 and 720, measuring performance via MSE. We progressively remove components from our full EntroPE architecture to create four configurations: (1) EntroPE (Full): Dynamic patching + Adaptive encoder + Fusion decoder; (2) EntroPE - Dynamic: Static patching + Adaptive encoder + Fusion decoder; (3) EntroPE - (Dynamic Patching + Adaptive Encoder): Static patching + Max pooling + Fusion decoder; (4) EntroPE - (Dynamic Patching + Adaptive Encoder + Fusion Decoder): Static patching + Max pooling + Flattened output.
Table 3: Component importance and dynamic vs. static patching comparison. Best and second-best are highlighted in red and blue.
Dataset T Full -EDP -EDP -APE -EDP -APE -FD
Dynamic Static(1) Static(8) Static(16) Pool + FD Pool + Flat
ETTh1 336 0.429 0.425 0.441 0.444 0.438 0.519
720 0.439 0.460 0.469 0.477 0.461 0.527
ETTh2 336 0.355 0.428 0.435 0.439 0.439 0.462
720 0.397 0.448 0.470 0.463 0.466 0.456
ETTm1 336 0.393 0.401 0.409 0.415 0.402 0.427
720 0.445 0.451 0.459 0.452 0.457 0.483
Weather 336 0.258 0.261 0.265 0.270 0.262 0.262
720 0.341 0.343 0.353 0.359 0.347 0.346
Dynamic vs. Static Patching. In addition to the component ablations, we compare EntroPE with dynamic patching against three static (fixed-length, non-overlapping) patching schemes. Experiments are conducted on ETTh1, ETTm1, and Weather datasets with a fixed input length of 96 and forecasting horizons of 336 and 720. Dynamic patching achieves the lowest MSE in most settings (Table 3), while single-token input (patch length of 1) remains competitive, notably without leveraging the Dynamic Patching scheme. However, the computational cost of this setting is substantially higher, as the self-attention mechanism must attend to every time step individually. Figure 3 illustrates the efficiency advantage of dynamic patching over static alternatives under different configurations.
Figure 3: Dynamic and Static(patch_length) patching with epoch duration.